# HidePrint: Hiding the Radio Fingerprint via Random Noise

### v1

@@ 4 Rationale of our Solution
where we have applied low-pass filtering to remove the component at frequency $2f_{c}$. We now assume that $\alpha(t)$ and $\phi(t)$ are affected by a synthetically generated noise with components $n_{\alpha}(t)$ and $n_{\phi}(t)$, respectively, as it follows:
@@
$\phi(t)$ $=$ $\phi^{\prime}(t)+n_{\phi}(t),$ (4)
where $\alpha^{\prime}(t)$ and $\phi^{\prime}(t)$ represent the actual amplitude and phase components of the baseband signal, respectively. We recall that $Q(t)=0$, when BPSK is considered. Under ideal conditions, Eq. 4 can be rewritten as:
@@
$\phi(t)$ $=$ $\{0,\pi\}+\epsilon_{\phi}(t)+n_{\phi}(t),$ (5)
where we considered the amplitude component $\alpha^{\prime}(t)=1+\epsilon_{\alpha}(t)$, i.e., a phasor (unitary amplitude) affected by the fingerprint $\epsilon_{\alpha}(t)$ and phase $\phi^{\prime}(t)=\{0,\pi\}+\epsilon_{\phi}(t)$, taking on the values either 00 or $\pi$ as a function of the transmitted bit, while $\epsilon_{\phi}(t)$ is the phase component.
We observe that the amplitude $n_{\alpha}(t)$ and phase $n_{\phi}(t)$ of the injected noise directly affect the components of the fingerprint, i.e., $\epsilon_{\alpha}(t)$ and $\epsilon_{\phi}(t)$. Therefore, the transmitter could hide such components by synthetically injecting noise into the baseband signal. Finally, we highlight that our analysis does not consider any noise coming from the channel, as well as multipath. We prove the reliability of our solution in the later sections of this paper, using real wireless measurements.
@@
As a toy example, consider Fig. 2, where we compare (i) the position of the IQ samples for the ideal BPSK modulation scheme (black crosses), (ii) an actual clear signal $r(t)$ (green area), and finally, (iii) the noise-obfuscated signal (red area), where the original signal $r(t)$ has been subjected to the noise components $n_{\alpha}(t)$ and $n_{\phi}(t)$. We highlight that the figure is the result of real communication in a wired link, where we applied a Gaussian noise of standard deviation 0.05 simultaneously on the real and imaginary parts of the signal ($10^{5}$ IQ samples). In the remainder of this work, we show that such noise intensity is enough to remove the radio fingerprint of the transmitter while negligibly affecting the quality of the signal, i.e., the Signal to Noise Ratio (SNR).
@@
Figure 2: A toy example of our HidePrint solution: the device fingerprint is removed from the clear signal (green area) by adding random noise, thus obtaining the signal dispersed in the red area (signal reshaped by HidePrint). The quality of the received signal remains almost the same.
5 Measurement setup
@@
Software and hardware setup. Our measurement setup is constituted by a set of Software Defined Radios, where the receiver is always an Ettus USRP X410 [31], and we considered a pool of ten (10) transmitters being the USRP B200-mini-i [32], as shown in Fig. 3. We consider a general-purpose laptop running Ubuntu 24.04 LTS running GnuRadio 3.10 (with Python 3.12) and the related flowcharts to control the receiver and each of the transmitters (one per time).
@@
Figure 3: Our measurement setup: 10 transmitters (USRP B200-mini-i) are connected (one per time) to the receiver (USRP X410). We collect data either via an Ethernet cable (USRP X410) or with a USB connection (USRP B200-mini-i). We consider two scenarios: a wired connection between the transmitter and the receiver (green) and a wireless link, where we deployed two Vert900 antennas (red). For both the cases, we considered an attenuator of 30dB.
@@
The flowchart for the transmitter (USRP B200-mini-i) involves a standard BPSK modulation scheme, where a repeating sequence of bytes spanning between 0 and 255 is generated considering the block Vector Source. The sequence is processed and converted to a stream of bits and, finally, to a stream of complex numbers where the noise $(n_{\alpha}(t),n_{\phi}(t))$ ( recall Sect. 4) is added before the Root Raised Cosine Filter. Finally, our transmitter chain features a USRP Sink block used to transmit the signal. The RRC filter performs an upsampling of 4 samples per symbol, i.e., from a symbol rate of 250K IQ symbols per second to $10^{6}$ samples per second. All measurements considered in this work assume a carrier frequency $f_{c}=900$MHz and a normalized transmitter power of 0.7.
The flowchart for the receiver (ran on the USRP X410) involves a USRP Source block with sample rate $10^{6}$ (same as the transmitter, according to the modified Nyquist theorem for complex signals [33]), carrier frequency $f_{c}=900$MHz, and normalized transmitter gain of 0.7, followed by a Costas Loop block, an adaptive gain control AGC block, a Root Raised Cosine Filter downsampling from 1M sample per second to 512k samples per seconds, and finally, a  Symbol Sync block implementing the Gardner algorithm for the time error detector and a polyphase filter bank as the interpolating resampler, downsampling the stream from 512K to (the original) 256K samples (symbols) per seconds.
We stress that the flowcharts considered at both the transmitter and the receiver are designed to reflect as much as possible the standard modulation and demodulation chains implemented in the vast majority of the commercial off-the-shelf transmitters and receivers. In fact, our flowcharts do not expose (or hide) features that might help or prevent the RFF process. As an example, the usage of the Costas Loop block—a standard component in any receiver—allows the recovery of the carrier and the synchronization between the clocks between the receiver and the transmitter, but unfortunately, it prevents (non-standard) solutions, like [23], which exploit CFO to perform radio fingerprinting.
Communication link. In this work, we consider two types of communication links, i.e., a wired link constituted by an RF cable and a wireless link, where the signal is transmitted and received by using two Vert900 antennas (as per Fig. 3). For both communication links, we consider a signal attenuator of 30dB between the transmitter and the receiver.
@@
The choice of both wired and wireless links aims to understand the impact of wireless channel propagation and multipath fading on the anonymization process. In fact, as introduced in Sect. 4, multipath fading represents another source of noise that potentially helps to prevent the identification of the transmitter at the receiver side.
@@
Dataset collection. We collected two datasets of measurements for a total of approximately 167 GB of data, available at [34]. The dataset collected from the wired link ($\approx$ 134GB) is constituted by 240 measurements, i.e., the result of 10 transmitters, 4 types of noise (no noise, Gaussian, Impulse, Laplacian, and Uniform), and 6 noise levels. In contrast, the dataset collected from the wireless link ($\approx$ 33GB) is constituted by 60 measures, i.e., the result of 10 transmitters, 1 types of noise (Gaussian), and 6 noise levels. Each measurement lasts for 300 seconds, thus providing 750M of IQ samples (recall that the considered symbol rate was 250k symbols per second).
6 Device fingeprinting
@@
A critical task in the RFF technique is to mitigate the impact of multipath fading in the fingerprinting process (either during training or testing). As discussed in Sec. 2, several techniques have been proposed during the years, including preprocessing of the IQ samples, adaptation of (only) the first layers of the neural network, up to designing from scratch neural networks able to handle IQ samples. Training and testing (directly) on IQ samples require adapting the neural network (at least for the initial layers) and an ad-hoc preprocessing technique to mitigate the impact of the multipath fading. Recent contributions, e.g.,  [8, 2, 28, 16, 9, 26, 27, 29], show the effectiveness of an orthogonal technique that involves preprocessing IQ samples into images and then using state-of-the-art DL techniques to discriminate and classify images. The root idea consists of generating images from IQ samples while mitigating the impact of the noise by averaging the spatial position of the IQ samples in the IQ plane. As CNNs are a well-known class of DL neural network specifically designed to classify images, image-based RFF techniques use CNNs to discriminate images generated from samples of a specific transmitter, thus performing device fingerprinting. In the following, we recall the most important details associated with the considered methodologies and procedures while pointing the reader to relevant literature for more detailed information.
@@
Data pre-processing. Our reference input is a matrix of 2 columns (In-phase and Quadrature components) and 750M of rows, one row for each IQ sample. As depicted in Fig. 2, the vast majority of the IQ plane is not used, and therefore we split the IQ plane into two parts (the positive and the negative In-phase components) and then cut out the 0.005 and 0.0095 quantiles of each axis. Considering the reference value of $10^{5}$ IQ samples per image (chunks extracted from the measurement), such a strategy involves removing a total of $2,000$ samples per chunk, allowing us to center the image on each cloud of IQ samples while removing outliers, as depicted in Fig. 4(a). The subsequent step consists of computing the bi-variate histogram on the data from Fig. 4(a), as depicted in Fig. 4(b). The number of bins of the bi-variate histogram is defined according to the size (width, height) of the images to be generated, and in turn, according to the input layer of the CNN. As discussed in the following, we consider the CNN ResNet-18, already implemented in Matlab2023b, thus setting the image size to 224-by-224 pixels. Figure 4(b) wraps up on the performed computations: the bi-variate histogram counts the IQ samples per bin (i.e., the density), and such value is considered as the value of the pixel in the generated image (lower part of Figure 4(b)). The number of IQ samples considered per image ($10^{5}$) should be carefully evaluated. Indeed, a high number of IQ samples per image might generate too many tiles of the bi-variate histogram characterized by values bigger than 255, which is a non-consistent pixel value, thus leading to significant information loss. We carefully estimated the value of the number of IQ samples per image, i.e., $10^{5}$, with empirical testing and in accordance with [9, 8].
@@
Figure 4: We split the data collected (IQ samples) into chunks of $10^{5}$ samples and organized them in one only cloud, i.e., the left cloud of the BPSK modulation is mirrored on the right side. We then compute a bi-variate histogram and consider the output as an image, ready to be processed by a DL algorithm.
Deep Learning. We consider two DL architectures, i.e., CNN and AE, as a function of the attack performed by the adversary (recall Sect. 3). When considering the adversary $\mathcal{A}_{1}$, we assume the adversary targets a finite pool of devices and has already trained a model on the full set. Each device in the pool is willing to keep its anonymity at maximum (k-anonimity, as per the definition in [35]), i.e., accuracy equal to 0.1, with $k=10$ devices in the pool. When considering $\mathcal{A}_{2}$, we consider a targeted attack, i.e., the adversary trains his model on a dataset coming from a specific device, and subsequently, he wants to detect the presence of such device in the radio spectrum. This problem can be addressed as a one-class classification problem, where the samples from the test set are detected as either legitimate, if they are recognized to be coming from the target device, or anomaly, if they are coming from any other device.
@@
Our reference CNN ResNet-18 is already implemented in Matlab R2023b and pre-trained on ImageNet [36], in accordance with the Transfer Learning design principle. The main objective of this work is not to find the best configuration for the RFF task, which has been addressed in previous work [8]. We choose ResNet-18 as it represents a good trade-off between accuracy and training time. Moreover, ResNet-18 cannot be used as is: its last layer requires adaptation to the number of classes (transmitters) in our problem, i.e., 10. Finally, we partially retrained the network by considering a stop criterion according to which the training is interrupted when the accuracy during the validation process achieves a variance across the last 5 iterations less than 0.5.
Our reference autoencoder is characterized by 64 neurons in the hidden layer, an encoder transfer function implemented with the logistic sigmoid function, a decoder transfer function implemented with a linear transfer function, a sparsity regularization coefficient of 0.5, and finally, a $L_{2}$ weight regularizer coefficient being equal to 0.01.
@@
While CNNs perform a prediction (class) for each provided image, the autoencoder outputs a reconstruction of the input. In fact, the autoencoder is constituted by an encoder and a decoder. During the training phase, the encoder learns a set of features (latent representation) and these features are used by the decoder to reconstruct the input. This makes the autoencoder applicable in scenarios where a prediction of inputs not previously seen is requested. Given the structure of the autoencoders, a fundamental metric is represented by the reconstruction error since it represents how the reconstructed output is “far away” from the input when reconstructed with the input’s features. In the remainder of this work, we will use the Mean Square Error (MSE) to assess if the image (and the associated IQ samples) belongs to the purported transmitter. An image with $MSE<\tau$ is predicted to be generated from the target transmitter, while an image with $MSE>\tau$ is predicted to be generated from another transmitter. For computing $\tau$, we adopt Eq. 6, as proposed by the authors in [37].
$\tau=E(MSE_{train})+3.5\cdot\sigma(MSE_{train}),$ (6)
where $E(\circ)$ and $\sigma(\circ)$ represent the mean and the standard deviation, respectively, of the MSE obtained during the validation process.
7 Results
@@
In this section, we present the results of our analysis related to two models of adversaries, i.e., $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ (recall Sect 3), while considering increasing prior knowledge. Indeed, we can make two different assumptions on the background knowledge of the adversary, independently of the model. We can assume that the adversary was able to train a model either on noise-free samples or on a combination of noise-free and noisy samples (all samples).
7.1 Wired Link - Adversary $\mathcal{A}_{1}$

### v2

@@ 4. Rationale of our Solution
where we applied low-pass filtering to remove the component at $f=2f_{c}$. We now assume that $\alpha(t)$ and $\phi(t)$ are affected by synthetically-generated noise with components $n_{\alpha}(t)$ and $n_{\phi}(t)$, i.e.:
@@
(4) $\phi(t)$ $=$ $\phi^{\prime}(t)+n_{\phi}(t),$
where $\alpha^{\prime}(t)$ and $\phi^{\prime}(t)$ represent the actual amplitude and phase components of the baseband signal, respectively. We recall that $Q(t)=0$, with BPSK. Under ideal conditions, Eq. 4 becomes:
@@
(5) $\phi(t)$ $=$ $\{0,\pi\}+\epsilon_{\phi}(t)+n_{\phi}(t),$
@@
where we considered the amplitude component $\alpha^{\prime}(t)=1+\epsilon_{\alpha}(t)$, i.e., a phasor (unitary amplitude) affected by fingerprint $\epsilon_{\alpha}(t)$ and phase $\phi^{\prime}(t)=\{0,\pi\}+\epsilon_{\phi}(t)$, taking on the values either 0 or $\pi$ as a function of the transmitted bit, while $\epsilon_{\phi}(t)$ is the phase component. Note that $n_{\alpha}(t)$ and $n_{\phi}(t)$ directly affect the components of the fingerprint $\epsilon_{\alpha}(t)$ and $\epsilon_{\phi}(t)$. Therefore, the transmitter could hide such components by synthetically injecting noise into the baseband signal. Finally, we highlight that our analysis does not consider any noise coming from the channel, as well as multipath. We prove the reliability of our solution in the later sections of this paper, using real wireless measurements. As a toy example, consider Fig. 2, where we compare: (i) the position of the IQ samples for the ideal BPSK modulation scheme (black crosses), (ii) an actual clear signal $r(t)$ (green area), and finally, (iii) the noise-obfuscated signal (red area), where the original signal $r(t)$ has been subjected to the noise components $n_{\alpha}(t)$ and $n_{\phi}(t)$. We highlight that the figure is the result of real communication in a wired link, where we applied a Gaussian noise of standard deviation 0.05 simultaneously on the real and imaginary parts of the signal ($10^{5}$ IQ samples). In the remainder of this work, we show that such noise intensity is enough to remove the radio fingerprint of the transmitter while negligibly affecting the quality of the signal, i.e., the Signal to Noise Ratio (SNR).
@@
Figure 2. A toy example of our HidePrint solution: the device fingerprint is removed from the clear signal (green area) by adding random noise, thus obtaining the signal dispersed in the red area (signal reshaped by HidePrint).
5. Measurement setup
@@
Software and hardware setup. Our measurement setup consists of a receiver Ettus USRP X410 (X410_datasheet), and ten (10) transmitters USRP B200-mini-i (b200_datasheet) (see Fig. 3). We consider a general-purpose laptop with Ubuntu 24.04 LTS running GNURadio 3.10 (with Python 3.12) and the related flowcharts to control the receiver and each of the transmitters (one per time).
@@
Figure 3. We use 10 transmitters (USRP B200-mini-i) connected (one per time) to the receiver (USRP X410). We consider two scenarios: a wired connection between the transmitter and the receiver (green) and a wireless link with two Vert900 antennas (red). For both, we use 30dB attenuation.
@@
The flowchart for the transmitter (USRP B200-mini-i) involves a standard BPSK modulation scheme, transmitting a repeating sequence of bytes spanning between 0 and 255, generated through the block Vector Source. The sequence is processed and converted to a stream of bits and, finally, to a stream of complex numbers where the noise $(n_{\alpha}(t),n_{\phi}(t))$ (recall Sect. 4) is added before the Root Raised Cosine Filter. Finally, we use a USRP Sink block to transmit the signal. The RRC filter performs an upsampling of 4 samples per symbol, i.e., from a symbol rate of 250K IQ symbols per second to $10^{6}$ samples per second. All our measurements consider a carrier frequency $f_{c}=900$ MHz and a normalized transmitter power of 0.7. The flowchart for the receiver involves a USRP Source block with sample rate $10^{6}$ (same as the transmitter, according to the modified Nyquist theorem for complex signals (sampling_same)), carrier frequency $f_{c}=900$ MHz, and normalized receiver gain of 0.7, followed by a Costas Loop block, an adaptive gain control AGC block, a Root Raised Cosine Filter downsampling from 1 M samples per second (sps) to 512k sps, and finally, a Symbol Sync block implementing the Gardner algorithm for the time error detector and a polyphase filter bank as the interpolating resampler, downsampling the stream from 512K sps to (the original) 256K sps. We remark that our flowcharts reflect as much as possible the standard modulation and demodulation chains implemented in commercial off-the-shelf devices. At the same time, our flowcharts do not expose (or hide) features that might help or prevent the RFF process. For example, the Costas Loop block—a standard component in any receiver—allows carrier recovery and TX-RX clocks synchronization, but it prevents (non-standard) solutions like (bechir_2024_tmlcn), which exploit CFO to perform RFF. Finally, we consider the SNR as our main control variable, so to make our results independent of the power levels and the geometry of the deployment.
@@
Communication link. We consider both a wired link (RF cable) and a wireless link with two Vert900 antennas (Fig. 3). We also use a signal attenuator of 30dB between the transmitter and the receiver, so as to understand the impact of wireless channel propagation and multipath fading on the anonymization process. In fact, as introduced in Sect. 4, multipath fading represents yet another source of noise potentially affecting RFF performance.
@@
Dataset collection. We collected two datasets for a total of approximately 167 GB of data, available at (dataset_anon). The artifacts uploaded on the submission platform consist of all the post-processed data, processing scripts and intermediate data to generate the figures. The raw data (167GB) are too large to be shared anonymously. Thus, they will be made public at a later moment. The dataset collected from the wired link ($\approx$ 134GB) includes 240 measurements, i.e., the result of 10 transmitters, 4 types of noise (no noise, Gaussian, Impulse, Laplacian, and Uniform), and 6 noise levels. The dataset collected from the wireless link ($\approx$ 33GB) includes 60 measures, i.e., the result of 10 transmitters, 1 types of noise (Gaussian), and 6 noise levels. Each measurement lasts for 300 seconds, thus providing 750M of IQ samples (sample rate 250k sps).
6. Device fingerprinting
@@
In this section, we present the RFF techniques considered for analysis. We recall that mitigating multipath fading is critical for improving RFF performance, and several techniques have been proposed, e.g., preprocessing of the IQ samples, adaptation of (only) the first layers of the neural network, up to designing from scratch neural networks able to handle IQ samples (see Sec. 2). Nevertheless, training and testing (directly) on IQ samples require adapting the neural network (at least for the initial layers) and an ad-hoc preprocessing technique to mitigate the impact of the multipath fading. Recent state-of-the-art contributions, e.g., PAST-AI (oligeri2023tifs), show the effectiveness of preprocessing IQ samples into images and then using state-of-the-art DL techniques to classify images. The root idea consists of generating images from IQ samples while mitigating the impact of the noise by averaging the spatial position of the IQ samples in the IQ plane. As CNNs are a well-known class of DL neural networks specifically designed to classify images, image-based RFF techniques use CNNs to discriminate images generated from samples of a specific transmitter, thus performing device fingerprinting. We recall below the most important steps associated with the considered methodologies. Interested readers can refer to the relevant literature for details.
@@
Data pre-processing. Our input is a matrix of 2 columns (IQ components) and 750M rows, one row for each IQ sample. As depicted in Fig. 2, the vast majority of the IQ plane is not used, and therefore we split the IQ plane into two parts (the positive and negative In-phase components) and then cut out the 0.005 and 0.0095 quantiles of each axis. Considering the reference value of $10^{5}$ IQ samples per image (chunks extracted from the measurement), such a strategy involves removing a total of $2,000$ samples per chunk, allowing us to center the image on each cloud of IQ samples while removing outliers (Fig. 4(a)). The subsequent step consists of computing the bi-variate histogram on the data from Fig. 4(a), as depicted in Fig. 4(b). The number of bins of the bi-variate histogram is defined according to the size (width, height) of the images to be generated, and in turn, according to the input layer of the CNN. As discussed in the following, we consider the CNN ResNet-18, already implemented in Matlab2023b, thus setting the image size to 224-by-224 pixels. Figure 4(b) wraps up on the performed computations: the bi-variate histogram counts the IQ samples per bin (i.e., the density), and such value is considered as the value of the pixel in the generated image (lower part of Figure 4(b)). The number of IQ samples considered per image ($10^{5}$) should be carefully evaluated. Indeed, a high number of IQ samples per image might generate too many tiles of the bi-variate histogram characterized by values bigger than 255, which is a non-consistent pixel value, thus leading to significant information loss. We carefully estimated the value of the number of IQ samples per image, i.e., $10^{5}$, with empirical testing and in accordance with (oligeri2023tifs; alhazbi2023_acsac).
@@
Figure 4. We split the data collected (IQ samples) into chunks of $10^{5}$ samples and organized them in one only cloud, i.e., the left cloud of the BPSK modulation is mirrored on the right side. We then compute a bi-variate histogram and consider the output as an image, ready to be processed by a DL algorithm.
@@
Deep Learning. We consider two DL architectures, i.e., CNN and AE, as a function of the attack performed by the adversary (recall Sect. 3). When considering the adversary $\mathcal{A}_{1}$, we assume the adversary targets a finite pool of devices and has already trained a model on the full set. Each device in the pool is willing to keep its anonymity at maximum (k-anonymity, as per the definition in (Samarati1998ProtectingPW)), i.e., accuracy equal to 0.1, with $k=10$ devices in the pool. When considering $\mathcal{A}_{2}$, we consider a targeted attack, i.e., the adversary trains his model on a dataset coming from a specific device, and subsequently, he wants to detect the presence of such a device in the radio spectrum. This problem can be addressed as a one-class classification problem, where the samples from the test set are detected as either legitimate, if they are recognized to be coming from the target device, or anomaly, if they are coming from any other device. Our reference CNN ResNet-18 is already implemented in Matlab R2023b and pre-trained on ImageNet (deng2009imagenet), in accordance with the Transfer Learning design principle. The main objective of this work is not to find the best configuration for the RFF task, which has been addressed in previous work (alhazbi2023_acsac). We choose ResNet-18 in line with (oligeri2023tifs), as it represents a good trade-off between accuracy and training time. Moreover, ResNet-18 cannot be used as it is: its last layer requires adaptation to the number of classes (transmitters) in our problem, i.e., 10. Finally, we partially retrained the network by considering a stop criterion according to which the training is interrupted when the accuracy during the validation process achieves a variance across the last 5 iterations less than 0.5.
@@
Our reference autoencoder, in line with (oligeri2023tifs), is characterized by 64 neurons in the hidden layer, the logistic sigmoid function as encoder transfer function, a decoder transfer function implemented with a linear transfer function, a sparsity regularization coefficient of 0.5, and finally, a $L_{2}$ weight regularizer coefficient equal to 0.01. The autoencoder is constituted by an encoder and a decoder, and it outputs a reconstruction of the input. During the training phase, the encoder learns a set of features (latent representation), which is then used by the decoder to reconstruct the input. This makes the autoencoder applicable in scenarios where a prediction of inputs not previously seen is requested. Given the structure of the autoencoders, a fundamental metric is represented by the reconstruction error since it represents how the reconstructed output is “far away” from the input when reconstructed with the input’s features. In the remainder of this work, we will use the Mean Square Error (MSE) to assess if the image (and the associated IQ samples) belongs to the purported transmitter. An image with $MSE<\tau$ is predicted to be generated from the target transmitter, while an image with $MSE>\tau$ is predicted to be generated from another transmitter. For computing $\tau$, we adopt the standard formula in Eq. 6, where $E(\circ)$ and $\sigma(\circ)$ represent the mean and the standard deviation, respectively, of the MSE obtained during the validation process.
(6) $\tau=E(MSE_{val})+3.5\cdot\sigma(MSE_{val}),$
7. Results
@@
In this section, we present the results of our analysis related to two models of adversaries, i.e., $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ (recall Sect. 3), while considering an increasing a priori knowledge of the transmitters. Indeed, we can make two different assumptions on the background knowledge of the adversary, independently of the model. We can assume that the adversary was able to train a model either on noise-free samples or on a combination of noise-free and noisy samples (all samples).
7.1. Wired Link - Adversary $\mathcal{A}_{1}$